As a student you are told to acquire the following textbooks; 5mathematics, 3physics, 8chemistry, 10biology, 12english, each of these costs N900, 700, 650, 500 and 400 respectively. Find the total cost of the textbooks?, find the sub-take cost of the textbooks?, find the VAT of the textbook if the textbook vat is 5%?, find the actual cost of the textbook?
1
Expert's answer
2019-07-28T17:35:47-0400

5mathematicsx90v0="5X900=4500"

3physicsx700="3X700=2100"

8chemistryx650="8X650=5200"

10biologyx500="10X500=5000"

Total cost =(4500+2100+5200+5000+4800)=21600

Sub-take cost= "21600\/2=10800"

Actual cost="21600X100\/105=20571.48"

VAT= "21600-20571.43=1028.57"